Find two consecutive integers such that three times the second integer is fourteen less than five times the first integer. Find the integers

Answer by ankor@dixie-net.com(22740)

Find two consecutive integers
n, (n+1)
such that three times the second integer is fourteen less than five times the first integer.
3(n+1) = 5n - 14
3n + 3 = 5n - 14
3 + 14 = 5n - 3n
17 = 2n
n = 17/2
n = 8.5; obviously something wrong here.
:
They do satisfy the equation but are not integers. A bad problem